OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8

OpenStudy (texaschic101):

x = -4y + 4....I am going to rearrange this one... x + 4y 4. x + 4y = 4 --> (-2) 2x + 8y = 8 -------------- -2x - 8y = -8 (result of multiplying by -2) 2x + 8y = 8 --------------add 0 = 0 Infinite solutions

OpenStudy (anonymous):

Thank you, I thought it was that, but I just wanted to make sure..

OpenStudy (texaschic101):

no problem....sometimes just getting a confirmation is good enough.

OpenStudy (goformit100):

First of all @hannahford1 : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct